write the first seven terms, the 20th term, and the sum of the first 20 terms for the arithmetic sequence where a =1 and d=2.5
1

Answer by mananth(16946)

a=1
d=2.5
t1=a =1
f2 = a+d = 3.5
t3 = a+2d = 4
t4 = a+3d = 5.5
t5 = a+4s = 7
t6 = a+5d = 8.5
t7 = a+6d = 10
tn = a+(n-1)d
t20 = 1+19*1.5
=28.5
Sum of n rems
sn = N/2[2a +(n-1)d]
S20 = 20/2 [2+19*1.5]
=305